OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -7).

hero (hero):

@lornbeach, what is the directrix of this parabola?

OpenStudy (anonymous):

I'm not sure, I have an equation but not sure how to apply it

hero (hero):

The focus and the directrix are always the same distance from the origin.

hero (hero):

So if the focus is (0 , -7), then the directrix would be y = 7. Expressed as a point it would be (x, 7). Once you have two points, the focus (0,-7) and the directrix (x, 7), you can insert them into the following formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\)

OpenStudy (anonymous):

Okay so what is x, x1, x2, y, y1, y2

OpenStudy (anonymous):

Like would x be x because we don't know that point

hero (hero):

When figuring out the equation of the parabola, you do the same thing. You insert the given points \(0, -7) and (x, 7) into that distance formula I gave you.

OpenStudy (anonymous):

Okay so it would be 7--7/x-0?

hero (hero):

I don't mean to confuse you. Let me delete that slope formula.

hero (hero):

I thought that maybe you would understand

hero (hero):

Insert those points in to THIS formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\)

OpenStudy (anonymous):

Ahh gotcha

hero (hero):

\((x_1, y_1) = (0, -7)\) \((x_2, y_2) = (x, 7)\)

OpenStudy (anonymous):

(x-0)^2+(y--7)^2=(x-x)^2+(y-7)^2

hero (hero):

Yes, that's mostly right, except for one thing. I would write it this way: (x-0)^2+(y-(-7))^2=(x-x)^2+(y-7)^2

hero (hero):

Make sure you emphasize -(-7) since it equals + 7 Putting two negatives next to each other doesn't mean anything without parentheses.

hero (hero):

Nevertheless from there you should be able to simplify and expand.

OpenStudy (anonymous):

Thank you(:

hero (hero):

Did you figure out the correct answer? If so you should post what you believe it is.

OpenStudy (anonymous):

I haven't done the work yet haha